# Batch learning equals online learning in Bayesian supervised learning

### v1

@@ 7. Final remarks
    In this paper we proved that batch Bayesian learning equals Bayesian online learning under the assumption of conditionally independent data, making Bayesian learning more efficient in the presence of complex data, and our Theorem 4.1 generalizes the celebrated Kalman filter.
@@
    Bayesian regression learning with corrupted measurement can be extended to nonlinear spaces $\mathcal{Y}$} where we can model measurement error using probability measures, e.g., for homogeneous Riemannian manifolds $\mathcal{Y}$}. Corollary 4.2 can be extended for Bayesian regression learning with corrupted measurement in the same way.
@@
    It is important to find a suitable concept of predictive consistency of Bayesian supervised learning which would agrees with the classical concept and the concept in a recent work by P. Koerpernik and F. Pfaff. A possible solution is to introduce the notion of predictive consistency at a finite subset $A\subset{\mathcal{X}}$. If ${\mathcal{X}}=1$ and the sampling operator is Markov kernel this concept is the notion of posterior consistency in classical Bayesian statistics [GV2017, §6.8.3].

### v2

@@ 7. Final remarks
    In this paper we proved that batch Bayesian learning equals Bayesian online learning under the assumption of conditionally independent data, making Bayesian learning more efficient in the presence of complex data. Our Theorem 4.1 generalizes the celebrated Kalman filter.
@@
    Bayesian regression learning with corrupted measurements can be extended to nonlinear spaces $\mathcal{Y}$} where we can model measurement error using probability measures, e.g., for homogeneous Riemannian manifolds $\mathcal{Y}$}. Corollary 4.2 can be extended for Bayesian regression learning with corrupted measurements in the same way.
@@
    It is important to find a suitable concept of predictive consistency of Bayesian supervised learning which would agrees with the classical concept and the concept in a recent work by P. Koerpernik and F. Pfaff. A possible solution is to introduce the notion of predictive consistency at a finite subset $A\subset{\mathcal{X}}$ and consider posterior predictive distributions of the form ${\mathcal{P}}_{T_{m}|S_{n},\mu}$ where $T_{m}]\subset A$ and $\Pi_{\mathcal{X}}(S_{n})]\in A$, moreover, $\Pi_{\mathcal{X}}(S_{n})$ visits each element of $A$ infinitely many times. If ${\mathcal{X}}=1$ and the sampling operator is Markov kernel this concept is the notion of posterior consistency in classical Bayesian statistics [GV2017, §6.8.3]. Furthermore, this notion also agrees with the condition of recurrent density entered in the concept of posterior consistency of Gaussian process regressions in [KP2021].